# Open Wireless Digital Twin: End-to-End 5G Mobility Emulation with OpenAirInterface and Sionna RT

### v3

@@ Abstract
This study presents an end-to-end wireless digital twin platform constructed using open-source software and open data to enhance the evaluation of mobile communication systems. The proposed open wireless digital twin (OWDT) integrates OpenAirInterface (OAI) for 5G NR protocol stack emulation and NVIDIA Sionna RT for high-resolution ray-tracing-based radio propagation modeling. This integration enables the realistic emulation of 5G wireless communication in mobility scenarios on a CPU-based Linux system, leveraging real-world building data to bridge the gap between theoretical simulations and real-world deployment. The platform also incorporates OAI FlexRIC, which is an implementation aligned with the O-RAN near-real-time RAN Intelligent Controller (near-RT RIC), to dynamically monitor the key performance indicators (KPIs). Through extensive evaluation in urban environments, this study demonstrated the validity of the emulation framework, revealing its capability to replicate real-world communication dynamics with high fidelity. The results underscore the potential of the OWDT to accelerate wireless system development, reduce experimental costs, and optimize network configurations.
@@
Digital twin, emulation, SDR, mobility, 5G NR, O-RAN, OpenAirInterface, Sionna RT, SUMO, PLATEAU.
1 INTRODUCTION
1.1 BACKGROUND
In the research and development (R&D) of wireless-assisted connected autonomous vehicles (CAVs) for future intelligent transportation systems (ITS), it is necessary to measure and evaluate changes in various mobile communication key performance indicators (KPIs) as vehicle-mounted antennas move [1, 2]. However, challenges arise owing to the difficulty in ensuring the reproducibility of the measurement data caused by the complex radio propagation characteristics of real-world physical environments. Additionally, the complexity of the standard specifications for mobile communications, such as 5G NR, poses challenges in interpreting and evaluating the obtained data. To address these challenges, emulation technology, which allows experiments to be conducted under ideal conditions using primarily real equipment, is gaining attention as a preliminary step before conducting experiments in increasingly complex field environments.
Next-generation wireless communication systems (e.g., 6G) typically undergo research, development, and demonstration stages before being deployed in society. In particular, it is important to evaluate the performance of a newly developed communication system through demonstration experiments based on actual-use scenarios. At the same time, the demonstration itself is a significant technical challenge in the R&D of wireless technologies for industrial deployment, often referred to as one of the‘valleys of death.’ The challenges associated with implementation and evaluation include the following: (1) Differences between the simulation models and real equipment, such as RF front ends. (2) Demonstration experiments require not only infrastructure costs and preparation time but, in some cases, the acquisition of radio station licenses, resulting in significant time and financial constraints that limit the ability to repeat comparative evaluations. (3) The complexity of wireless communication system protocols increases as they become more advanced. (4) It is difficult to accurately simulate the radio-propagation environment when pre-evaluating actual demonstration scenarios.
@@ 1.1 BACKGROUND
Against this backdrop, solutions addressing challenges (1) and (2) have led to a growing demand beyond the wireless communication field for conducting scenario testing in virtual environments of digital twins that replicate the real world with high fidelity before conducting field experiments [3, 4]. Digital twin technology has been actively discussed within some standardization projects such as 3GPP and Open Radio Access Network (O-RAN), and is recognized as a promising tool for generating training data for artificial intelligence and machine learning (AI/ML) models used in the RAN and network control [5, 6, 7]. Training AI/ML models typically requires large volumes of diverse data; however, the amount and variety of data that can be collected from real mobile networks are inherently limited. By leveraging digital twins, it is possible to generate virtually unlimited synthetic data and execute scenarios that deliberately introduce network faults. This enables the collection of rich and diverse datasets that are otherwise difficult or impossible to obtain in real-world environments. The key technical aspects of realizing a digital twin include emulation technology, which ensures high fidelity and real-time responsiveness in the virtual space, as well as consistency with measurement data that link the virtual and real worlds. These features enable the verification of newly proposed methods before field demonstration experiments, the reproducibility of time-varying KPIs that are difficult to achieve in real-world communication systems, and the optimization of deployment parameters. Thus, the development of digital twin platforms and emulation-based validation is emerging as a promising methodology for bridging the technological gap between theoretical development, simulations, and demonstration experiments, thereby accelerating the cycle that connects research innovations to societal deployment. The development of digital twin platforms for wireless communications has been increasing [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18].
@@
This compares only 5G OSS projects capable of independently implementing and experimenting with each 5G function without relying on external simulators. ‘OTA’ denotes whether the framework can connect an RU and conduct OTA experiments; ‘Prop.’ indicates the presence of stochastic or deterministic channel model implementations; ‘O-RAN’ signifies that at least one component of the O-RAN architecture is implemented. |
@@
Regarding the challenge of (3), recent trends in wireless communication R&D have shown significant advancements in the development of various open-source software (OSS) and the use of open data. One of the key factors behind this global open innovation trend is the increasing sophistication, diversification, and complication of wireless communication technologies, which prevent a single company or organization from conducting full-stack research and development independently. Many commercial wireless communication technologies are standardized by organizations, which means that common global specifications are defined. Consequently, implementing these common specifications as OSS is an economically rational approach on a global scale. It is supported by engineers worldwide, regardless of any specific for-profit organization. From a user perspective, OSS allows the low-cost construction of research and development environments while also enabling flexible customizations that are difficult to achieve with commercial software packages. As shown in Table 1, 5G OSS implementations have emerged to perform functions related to the RAN, 5G Core Network (CN), User Equipment (UE), and O-RAN. Among them, OpenAirInterface (OAI) stands out as a widely used 5G OSS in research and development, offering detailed implementations across 4G LTE/5G NR CN, RAN, UE, and O-RAN, along with support for over-the-air (OTA) experiments involving software-defined radio (SDR) devices and commercial off-the-shelf (COTS) Open RAN Radio Units (O-RUs) [19, 20]. Free5GC, Open5GS, and Magma constitute three of the most prominent OSS implementations of 5G CN [21, 22, 23]. AETHER, led by the Open Networking Foundation, is a project that unifies the subprojects SD-Core, SD-RAN, and an O-RAN near-real-time RAN Intelligent Controller (near-RT RIC) built on µONOS [24]. The srsRAN project, including srsUE, delivers standalone RAN and UE functionality and enables OTA experimentation with SDRs or O-RUs [25]. UERANSIM is a software simulator that provides RAN and UE functionality for testing [26]. NVIDIA Sionna is a GPU-accelerated link-level simulator for wireless communication systems [27]. In addition, Sionna RT offers differentiable ray tracing for deterministic radio propagation simulations [28]. The O-RAN Software Community (O-RAN SC) supplies multiple components of the O-RAN architecture, including O-CUs, O-DUs, RICs, and an SMO framework [29]. 5G-LENA and Simu5G are discrete-event network simulators built on the ns-3 and OMNeT++ platforms, respectively [30, 31]. A more detailed comparison of OSS projects from an O‑RAN‑focused perspective has been reported in a survey by another group [32].
@@
The modeling of radio propagation, which falls under difficulty (4), has been studied for improved accuracy using both stochastic and deterministic approaches. The stochastic approach allows the construction of channel models that statistically approximate experimental measurement results by parameter fitting. However, when 3D building models and terrain data are available, deterministic approaches, such as ray tracing, can explicitly track propagation paths and perform high-fidelity channel calculations from a microscopic perspective. In particular, ray tracing excels at simulating the delay profiles in environments with significant multipath propagation. While multipath signals are treated as fading effects at the receiver, a trade-off exists between the number of delay taps of the signals and the computational processing cost. Real-time processing is essential for ensuring the emulation functionality of digital twins. Therefore, efficient computation mechanisms are required to maintain the necessary number of delay taps to reproduce realistic fading effects while ensuring the real-time calculation of received signals. This study focuses on performing real-time convolution processing on baseband signals using a precomputed time-evolving radio propagation channel in a CPU-only environment.
1.2 MOTIVATION
Based on the above background, this study proposes the integration of OAI and Sionna RT to construct an OSS-based end-to-end (E2E) open wireless digital twin (OWDT) platform. Here, E2E refers to the communication section from 5G CN through RAN, including gNB, to UE. Specifically, we perform wireless communication emulation that considers realistic propagation environments based on site-specific mobility scenarios using actual building shapes and geographic data while utilizing O-RAN functionalities to monitor KPIs in near-RT [33]. In general, digital twin-based communication involves replicating real-world communication environments within a virtual space [11, 9, 14]. This replica includes 3D building structures, material properties, and detailed information on gNBs and UE antennas, such as their positions, orientations, and movements. The higher the accuracy of this information, the more precisely the emulation can capture the real wireless communications. Some companies have introduced network digital twin (NDT) solutions in the commercial sector, incorporating proprietary AI technologies [34, 35], where the term ‘NDT’ is defined as ‘virtual replica of mobile network or part of one, that captures its attributes, behaviour and interactions’ noting that the mobile network includes both RAN and CN in 3GPP [5]. Recently, a full-stack RAN simulation environment that integrates ns-3 and Sionna RT, considering deterministic channel models, was reported [17, 18]. Although ns-3 is well suited for scalable network performance simulation, it abstracts the physical (PHY) layer signal processing and lacks a sub-microsecond discrete-time resolution. Therefore, there are inherent limits to capturing the time-domain effects of delayed multipath components, even when ray tracing is applied to model radio propagation. Open AI Cellular (OAIC) is an NSF‑supported OSS platform that stitches together srsRAN for RAN/UE operation and O‑RAN SC components (Near‑RT/Non‑RT RIC and supporting services), linking them to create an E2E, AI‑ready 5G/6G research testbed [36]. Because the OAIC supports SDR operation and performs channel emulation via Zero Message Queue (ZeroMQ), implementing an efficient channel convolution in baseband IQ signal processing should, in principle, enable time-varying channel emulation. However, because the OAIC does not yet support 5G CN functions, the ‘E2E’ emulation required in this study cannot be achieved. The Colosseum, a large-scale wireless network emulator developed at Northeastern University, implements an O-RAN-compliant digital twin environment [11]. Its channel emulation is performed in the Massive Channel Emulator (MCHEM), a cluster of FPGAs that convolves channel impulse responses (CIRs) with the baseband signals exchanged among many transmitter and receiver nodes. The convolution is executed by a 512‑tap FIR filter in which 4 arbitrary taps are set to nonzero values, a design that ensures high processing speed, lightweight data handling, and scalability with respect to the number of radio nodes. This 4‑tap approximation, however, limits the physical fidelity of each radio channel and forces the omission of bandwidth‑limiting filtering that should, in principle, be applied to the CIR before convolution.
Our work in this paper is based on OAI, which supports full-stack 5G NR without abstracting any layer and enables real-time operation on actual radio devices to explore the boundary between the PHY layer and radio propagation. This allows us to directly incorporate multipath fading effects into IQ-sample-level baseband signals. Since the radio processing runs in real time on real devices with high fidelity, the communications executed on our system can be described as an ‘emulation.’ Our proposed OWDT platform is built entirely on OSSs and requires no additional accelerators such as GPUs or FPGAs: a single, COTS CPU server is sufficient to perform real-time E2E 5G emulation between a gNB and a UE, without abstracting the PHY layer processing. This design is fully consistent with our goal of democratizing the wireless digital twin, i.e., making it open and accessible.
1.3 CONTRIBUTIONS AND RESTRICTIONS
The contributions of this study are as follows: 1) Democratization of Wireless Digital Twin through OSS- and CPU-Based Platform: The OWDT architecture is built from a standard CPU‑based Linux server, a low‑cost SDR device, a COTS UE module, and a laptop to control the UE, resulting in an extremely cost‑efficient setup. The workflow of this system is entirely composed of OSS or open data, such as OAI, Sionna RT, SUMO, OpenStreetMap, and PLATEAU, contributing to the democratization of the wireless digital twin environment. 2) Full Stack 5G Evaluation Environment with a lightweight near-RT RIC: We have built a full-stack 5G NR performance evaluation environment that integrates deterministic, non-abstracted radio propagation channel responses with FlexRIC, which is a lightweight, O-RAN-aligned near-RT RIC. 3) Near-RT E2E Emulation Testbed: We realized E2E communication emulation based on high-fidelity mobility scenarios in space and time by performing ‘real-time’ convolution operations on baseband signals using an SDR and pre-computed CIRs with monitoring KPIs through the xApp in ‘near-RT’. Real‑time convolution using up to 28 channel taps is achieved, improving over prior related works [11, 9]. 4) The CIR convolution implementation for OAI developed in this ‘Open Wireless Digital Twin’ project, along with a sample scenario, has been contributed as OAI source code and made publicly available as OSS [37].
This study is useful for gaining insights into the entire wireless communication system, where it is essential to integrate considerations of radio propagation as a physical phenomenon with digital signal processing. However, it is also important to acknowledge the following limitations: a) The OWDT environment does not account for channels varying on timescales shorter than the Doppler period; b) Although the contribution 3) has been fulfilled, the convolution incorporates only the 28 most dominant CIR taps, representing fewer than 20 % of the 146 taps computed in total; and c) Calibration with actual measured data has not been performed yet. Note that the current OWDT performs approximate emulation under these constraints. d) Faithfully capturing the impact of radio propagation on baseband signals comes at the expense of scalability: the number of radio nodes that can be supported is largely restricted.
1.4 PAPER ORGANIZATION
@@
The remainder of this paper is organized as follows. Section 2 introduces the E2E system architecture of the OWDT, incorporating the 5G NR protocol stack implemented by OAI, FlexRIC enabling O-RAN-aligned key performance measurement (KPM), and propagation channels generated by Sionna RT. It also describes the hardware environment in which these components are deployed. Next, the workflow for executing mobility emulation scenarios, such as vehicular driving, on the OWDT platform is explained. This section also introduces the OSSs and open data utilized in the workflow. Furthermore, it provides a detailed discussion of the convolution method used to apply CIRs to baseband signals, which serves as the integration point between OAI and Sionna RT. Section 3 evaluates the data obtained through ray tracing and the KPIs measured from the OAI after applying these data to the system while executing mobility scenarios on the platform. Finally, Section 4 concludes the paper with a discussion of the future perspectives.
2 METHODS

### v4

# Open Wireless Digital Twin: End-to-End 5G Mobility Emulation with OpenAirInterface and Ray Tracing

@@ Abstract
This study presents an end-to-end wireless digital twin platform constructed using open-source software and open data to enhance the evaluation of mobile communication systems. The proposed open wireless digital twin (OWDT) integrates OpenAirInterface (OAI) for Fifth-Generation New Radio (5G NR) protocol stack emulation and NVIDIA Sionna RT for high-resolution ray-tracing-based radio propagation modeling. This integration enables the realistic emulation of 5G wireless communication in mobility scenarios on a CPU-based Linux system, leveraging real-world building data to bridge the gap between theoretical simulations and real-world deployments. The platform also incorporates OAI FlexRIC, which is an implementation aligned with the O-RAN near-real-time RAN Intelligent Controller (near-RT RIC), to dynamically monitor key performance indicators (KPIs). Through extensive evaluation in urban environments, this study demonstrated the validity of the emulation framework, revealing its capability to replicate real-world communication dynamics with high fidelity. The results underscore the potential of the OWDT to accelerate wireless system development, reduce experimental costs, and optimize network configurations.
Index Terms:
@@
Digital twin, emulation, SDR, mobility, 5G NR, O-RAN, OpenAirInterface, Sionna RT, SUMO, PLATEAU.=-21pt
I Introduction
I-A BACKGROUND
In the research and development (R&D) of wireless-assisted connected autonomous vehicles (CAVs) for future intelligent transportation systems (ITS), it is important to evaluate dynamic aspects of key performance indicators (KPIs) through demonstration experiments based on real-world usage scenarios [1, 2]. Simultaneously, the demo constitutes one of the technical hurdles in wireless R&D aimed at industrial applications. The challenges in such demo experiments include (1) discrepancies that arise between simulation models and real devices, (2) the need in some cases to obtain radio station licenses, which increases time and financial costs, (3) growing protocol complexity as systems become more advanced, and (4) the difficulty of accurately reproducing the radio-propagation environment during pre-evaluation of the demo scenarios.
@@ I-A BACKGROUND
To address challenges (1) to (4), an approach has been proposed in which scenario testing is conducted in a digital twin virtual environment that replicates the real world with high fidelity, enabling evaluation prior to field experiments [3, 4]. Digital twins are being actively discussed in standardization projects, such as 3rd Generation Partnership Project (3GPP) and Open Radio Access Network (O-RAN). They not only enable the verification of new methods in virtual space prior to field demonstration experiments but also make it possible to reproduce time-varying KPIs that are difficult to observe in real communication systems and to optimize deployment parameters. Consequently, the development of digital twin platforms and their use for validation are attracting attention as promising methodologies for bridging the technical gaps among theory, simulation, and demonstration experiments, and for accelerating the cycle of translating research outcomes into societal deployment [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Furthermore, digital twins are regarded as a promising means of generating virtually unlimited training data for Artificial Intelligence / Machine Learning (AI/ML) models used for Radio Access Network (RAN) and network control [16, 17, 18]. Key technological elements for realizing digital twins include emulation technologies that guarantee high fidelity and real-time performance in virtual space, as well as consistency with measurement data that link the virtual and real worlds.
@@
| This compares only 5G OSS projects capable of independently implementing and experimenting with each 5G function without relying on external simulators. ‘OTA’ denotes whether the framework can connect an RU and conduct OTA experiments; ‘Prop.’ indicates the presence of stochastic or deterministic channel model implementations; ‘O-RAN’ signifies that at least one component of the O-RAN architecture is implemented. | ||||||
@@
In particular, regarding challenge (3), recent R&D in wireless communications has shown significant advances in the development of various open-source software (OSS) and the use of open data. Driving this open innovation are the increasing sophistication, diversification, and complexity of wireless communication technologies; in fact, it has become difficult for a single organization to complete full-stack R&D on its own. Many wireless technologies are standardized globally by standards bodies. Therefore, implementing these common specifications as OSS is an economically rational approach on a global scale. From a user perspective, OSS allows the low-cost construction of R&D environments while also enabling flexible customizations that are difficult to achieve with commercial software packages. As shown in Table I, Fifth-Generation New Radio (5G NR) OSS implementations have emerged to perform functions related to the RAN, 5G Core Network (CN), User Equipment (UE), and O-RAN. Among them, OpenAirInterface (OAI) stands out as a widely used 5G OSS in R&D, offering detailed implementations across 4G Long Term Evolution (LTE)/5G NR CN, RAN, UE, and O-RAN, along with support for over-the-air (OTA) experiments involving software-defined radio (SDR) devices and commercial off-the-shelf (COTS) O-RAN Radio Units (O-RUs) [19, 20]. Free5GC, Open5GS, and Magma constitute three of the most prominent OSS implementations of the 5G CN [21, 22, 23]. AETHER, led by the Open Networking Foundation, is a project that unifies the subprojects SD-Core, SD-RAN, and an O-RAN near-real-time RAN Intelligent Controller (near-RT RIC) built on µONOS [24]. The srsRAN project, including srsUE, delivers standalone RAN and UE functionality and enables OTA experimentation with SDRs or O-RUs [25]. UERANSIM is a software simulator that provides RAN and UE functionality for testing [26]. NVIDIA Sionna is a GPU-accelerated link-level simulator for wireless communication systems [27]. In addition, Sionna RT offers differentiable ray tracing for deterministic radio propagation simulations [28]. The O-RAN Software Community (O-RAN SC) supplies multiple components of the O-RAN architecture, including O-CUs, O-DUs, RICs, and an SMO framework [29]. 5G-LENA and Simu5G are discrete-event network simulators built on the ns-3 and OMNeT++ platforms, respectively [30, 31]. A more detailed comparison of OSS projects from an O-RAN-focused perspective has been reported in a survey by another group [32].
@@
The modeling of radio propagation, which falls under challenge (4), has been studied for improved accuracy using both stochastic and deterministic approaches. The stochastic approach allows the construction of channel models that statistically approximate experimental measurement results through parameter fitting. However, when 3D building models and terrain data are available, deterministic approaches, such as ray tracing, can explicitly track propagation paths and perform high-fidelity channel calculations from a microscopic perspective. In particular, ray tracing excels at simulating delay profiles in environments with significant multipath propagation. While multipath signals are treated as fading effects at the receiver, a trade-off exists between the number of delay taps of the signals and the computational processing cost. Real-time processing is essential for ensuring the emulation of digital twins. Therefore, efficient computation mechanisms are required to maintain the necessary number of delay taps to reproduce realistic fading effects while ensuring the real-time calculation of received signals.
I-B RELATED WORKS
In the commercial sector, some companies have introduced network digital twin (NDT) solutions that incorporate proprietary AI technologies [33, 34]. Here, ‘NDT’ is defined as ‘virtual replica of mobile network or part of one, that captures its attributes, behavior and interactions,’ noting that the mobile network includes both RAN and CN in 3GPP [16].
A full-stack RAN simulation environment that integrates ns-3 and Sionna RT, considering deterministic channel models, was reported [14, 15]. Although ns-3 is well-suited for scalable network performance simulation, it abstracts the physical (PHY) layer signal processing and lacks sub-microsecond discrete-time resolution. Therefore, there are inherent limits to capturing the time-domain effects of delayed multipath components, even when ray tracing is applied to model radio propagation.
Open AI Cellular (OAIC) is an NSF-supported OSS platform that stitches together srsRAN for RAN/UE operation and O-RAN SC components (near-RT/non-RT RIC and supporting services) to build an end-to-end (E2E), AI-ready 5G/6G research testbed [35]. The OAIC supports the SDR operation and performs channel emulation via Zero Message Queue (ZeroMQ). In principle, if an efficient channel convolution is implemented for baseband In-phase and Quadrature (IQ) signal processing, it should be possible to emulate time-varying channels. However, the OAIC currently lacks 5G CN functionality and thus, the ‘E2E’ emulation required in this study cannot be achieved.
Colosseum, a large-scale wireless network emulator developed at Northeastern University, implements an O-RAN-compliant digital twin environment [8]. Its channel emulation is performed in the Massive Channel Emulator (MCHEM), a Field-Programmable Gate Array (FPGA) cluster that convolves channel impulse responses (CIRs) with baseband signals exchanged among numerous transmitter and receiver nodes. The convolution uses a 512‑tap finite impulse response (FIR) filter in which 4 arbitrary taps are set to nonzero values, a design that ensures high processing speed, lightweight data handling, and scalability with respect to the number of radio nodes. This 4‑tap approximation, however, limits the physical fidelity of each radio channel and forces the omission of bandwidth-limiting filtering that should, in principle, be applied to the CIR before convolution.
Another closely related effort has recently emerged within OAI [36]. Implemented on Godot, it enables channel generation via scenario control over a representational state transfer (REST) application programming interface (API) using the functions provided by Sionna [37]. With CIRs produced by this implementation, it is also possible to simulate 5G communications while accounting for dynamic channel updates via OAI vrtsim, a radio simulator with a fixed timescale. Our implementation differs in that it runs on a best-effort basis on the CPU in the simulation mode and operates only in real time in the SDR mode.
I-C CONTRIBUTIONS AND LIMITATIONS
We summarize the contributions of the proposed Open Wireless Digital Twin (OWDT) architecture as follows.
1.
An O-RAN-aligned 5G E2E evaluation environment. We provide an evaluation environment positioned as shown in the bottom row of Table I that integrates the strengths of OAI and Sionna RT to realize full-stack, non-abstracted 5G NR mobile communications in software. Here, ‘E2E’ scope spans 5G CN, RAN, and UE. Furthermore, by leveraging OAI FlexRIC, KPIs can be monitored in near-RT [38].
2.
A fully OSS-based wireless digital twin. All software and data used to build the environment were open-source and open data. In addition, the CIR convolution implemented on OAI in this study and the sample scenarios were released as the OAI branch [39].
- 3.
4.
A low-cost, compact testbed. As a testbed, the OWDT consists of a standard Linux server that runs on the CPU, a low-end SDR, and a COTS UE module, enabling the realization of a low-cost experimental platform.
These contributions enable controlled yet realistic investigations of the PHY–propagation boundary within a laboratory setup, thereby promoting the democratization of wireless digital twin experimentation.
Meanwhile, the present OWDT has the following limitations, so the emulation remains approximate: (a) It does not model channels varying on time scales shorter than the Doppler period; (b) out of all the computed CIR taps, it uses only the dominant 28 taps. Although this is an expansion by several times over prior works, it remains a constraint; (c) it has not yet been calibrated against measured data; (d) prioritizing PHY fidelity comes at the expense of scalability, limiting the number of radio nodes that can be supported simultaneously.
I-D PAPER ORGANIZATION
@@
The remainder of this paper is organized as follows. Section II introduces the E2E system architecture of the OWDT, incorporating the 5G NR protocol stack implemented by OAI, FlexRIC enabling O-RAN-aligned key performance measurement (KPM), and propagation channels generated by Sionna RT. It also describes the hardware environment in which these components were deployed. Next, the workflow for executing mobility emulation scenarios, such as vehicular driving, on the OWDT platform is explained. This section also introduces the OSSs and open data utilized in the workflow. Furthermore, it provides a detailed discussion of the convolution method used to apply CIRs to baseband signals, which serves as the integration point between OAI and Sionna RT. Section III evaluates the data obtained through ray tracing and the KPIs measured from the OAI after applying these data to the system while executing the mobility scenarios on the platform. Finally, Section IV concludes the paper with a discussion of future perspectives. Table II summarizes the abbreviations used throughout the paper; all terms are defined at first use.
| Acronym | Expanded form | Acronym | Expanded form |
| 3GPP | 3rd Generation Partnership Project | OAI | OpenAirInterface |
| 5G NR | Fifth-Generation New Radio | OAIC | Open AI Cellular |
| 5GCN | 5G Core Network | O-CU | O-RAN Central Unit |
| AGC | Automatic Gain Control | O-DU | O-RAN Distributed Unit |
| AI | Artificial Intelligence | O-RAN | Open Radio Access Network |
| AI/ML | Artificial Intelligence / Machine Learning | O-RAN SC | O-RAN Software Community |
| API | Application Programming Interface | O-RU | O-RAN Radio Unit |
| BLER | Block Error Rate | OFDM | Orthogonal Frequency Division Multiplexing |
| BLIS | BLAS-like Library Instantiation Software | OMNeT++ | Objective Modular Network Testbed in C++ |
| CAVs | Connected and Autonomous Vehicles | ONOS | Open Network Operating System |
| CDL | Clustered Delay Line | OSM | OpenStreetMap |
| CIR | Channel Impulse Response | OSS | Open-Source Software |
| CN | Core Network | OTA | Over-the-Air |
| COTS | Commercial Off-The-Shelf | OWDT | Open Wireless Digital Twin |
| CPU | Central Processing Unit | PDP | Power Delay Profile |
| CP | Cyclic Prefix | PHY | Physical (layer) |
| DL | Downlink | PLY | Polygon (file format) |
| E2 | E2 Interface | PRB | Physical Resource Block |
| E2E | End-to-End | R&D | Research and Development |
| E2SM | E2 Service Model | RAN | Radio Access Network |
| FFT | Fast Fourier Transform | REST | Representational State Transfer |
| FIR | Finite Impulse Response | RF | Radio Frequency |
| FPGA | Field-Programmable Gate Array | RIC | RAN Intelligent Controller |
| FR2 | Frequency Range 2 | RMS | Root Mean Square |
| GI | Guard Interval | RSRP | Reference Signal Received Power |
| GPU | Graphics Processing Unit | RT | Ray Tracing |
| gNB | next-generation NodeB (5G base station) | SA | Standalone |
| IQ | In-phase and Quadrature | SDK | Software Development Kit |
| ISI | Inter-Symbol Interference | SDR | Software-Defined Radio |
| ITS | Intelligent Transportation Systems | SISO | Single-Input Single-Output |
| KPI | Key Performance Indicator | SRS | Sounding Reference Signal |
| KPM | Key Performance Measurement (O-RAN E2SM-KPM) | SUMO | Simulation of Urban MObility |
| LoS | Line-of-Sight | TDD | Time Division Duplex |
| LOD | Level of Detail | TDL | Tapped Delay Line |
| LTE | Long-Term Evolution | TR | Technical Report |
| LTS | Long-Term Support | TS | Technical Specification |
| MCHEM | Massive Channel Emulator | UE | User Equipment |
| MCS | Modulation and Coding Scheme | UL | Uplink |
| MIMO | Multiple-Input Multiple-Output | USRP | Universal Software Radio Peripheral |
| MLIT | Ministry of Land, Infrastructure, Transport and Tourism (Japan) | WSS | Wide-Sense Stationary |
| NDT | Network Digital Twin | xApp | O-RAN near-RT RIC Application |
| NR | New Radio | XML | Extensible Markup Language |
| ns-3 | network simulator 3 | ZeroMQ | Zero Message Queue |
II METHODS